Installation of the new BlueFors xLD fridge at UNSW in January 2017.
Note this page does not contain technical photos / information. That is elsewhere.

Monday 23rd - unpacking crates and installing frames

Today, with a huge team effort, the lab went from this to this (click for bigger images):

P1130752-med.jpg to P1130754-med.jpg
IMG_5702_800x600.jpg Low res timelapse video (9Mb): Day_1_900x600_crf_35.mp4

Tuesday 24th - pumping lines, gas cabinet

Today we hooked up the pumping lines, installed the top-loading mechanism, and connected the gas cabinet. We also took the cans off. The system went from this:

EM230283-med.jpg to EM240312-med.jpg and finally EM240367-med.jpg
Camera 1  
Camera 2 Low Resolution (5Mb) CanonDay2-900x600.mp4

Wednesday 25th - connecting power, water and gas.

Today we got 3 phase and single phase power and UPS connections made, while Oleh also managed to connect the nitrogen and cooling water systems. We also set up the additional vibration isolation on the Cryomech cold head, as well as installing the probe loading gallows.

EM250388.JPG EM250381.JPG EM250391.JPG

Thursday 26th - leak testing and checking the top-loading mechanism.

EM260398.JPG EM260501.JPG EM260408.JPG

Checking it goes in straight:

EM260497.JPG EM260489.JPG EM260521.JPG

Friday 27th - putting the magnet on, sealing up cans, pumping out and starting the cooldown.

Today would have been a shorter day if we hadn't discovered two leaks in the vacuum line and OVC cans.

EM260220.JPG EM260234.JPG EM270262.JPG IMG_5714.JPG

Mon 30th - cooling to base.

Today we enjoyed finding out that the magnet power supply needs two 10A mains sockets, and that the door to the room with our electrical switchboard needs a key that nobody on campus has. After considerable stuffing around Oleh and Juho loaded the mixture from an old dilution fridge (thanks Bob!), and we started cooling the system to base.

Is the 30 year old mixture clean or will it block the fridge? Will there be enough He3? Will it phase separate?


Tue 31st - optimising mixture.

We added 3l more He3 today, with help from Rod and Dave.
Before adding the rate of cooling at 10:40 wasn't great. Adding 2l more He3 and the rate of cooling at 13:00 was much better. After some cooling power tests we added a further litre of He3. System now goes below 7mK (15mK on the probe).


Wed 1st - Testing the magnets.

The system has a 5/1/1 vector magnet. The 5 and 1T magnets worked perfectly. The 1T magnet had a quench at 0.75T, and now doesn't pass current. So we will have to fix that. Until then we have a two axis vector magnet.

Thu 2nd - Changing probes.

Today we tested the probe removal and reinsertion process - all looks good.

Fri 3rd - More tests and warming up again.

Probe 2 reached 12mK, which is nice. After some more cooling power tests, we had to start warming the system up so that we can look at the magnet problem.

Sun 5th - Opening the system up again.

With a bit of encouragement, the magnet is now back at room temperature. Oleh and Juho opened the system up again to diagnose the magnet problem. Hopefully we can get a replacement part and fix it quickly.

Access Control

This page is meant to be visible by the whole world.


create new tag


Topic attachments
I Attachment Action Size Date Who Comment
Base1.PNGPNG Base1.PNG manage 75 K 30 Jan 2017 - 08:05 AlexHamilton  
CanonDay2-900x600.mp4mp4 CanonDay2-900x600.mp4 manage 5 MB 26 Jan 2017 - 15:11 AlexHamilton  
Day_1_900x600_crf_35.mp4mp4 Day_1_900x600_crf_35.mp4 manage 8 MB 24 Jan 2017 - 01:32 AlexHamilton  
EM230283-med.jpgjpg EM230283-med.jpg manage 293 K 26 Jan 2017 - 14:07 AlexHamilton  
EM240312-med.jpgjpg EM240312-med.jpg manage 255 K 26 Jan 2017 - 14:14 AlexHamilton  
EM240367-med.jpgjpg EM240367-med.jpg manage 236 K 26 Jan 2017 - 14:19 AlexHamilton  
EM250381.JPGJPG EM250381.JPG manage 195 K 29 Jan 2017 - 13:56 AlexHamilton  
EM250388.JPGJPG EM250388.JPG manage 139 K 29 Jan 2017 - 13:55 AlexHamilton  
EM250391.JPGJPG EM250391.JPG manage 185 K 29 Jan 2017 - 13:56 AlexHamilton  
EM260220.JPGJPG EM260220.JPG manage 163 K 29 Jan 2017 - 14:42 AlexHamilton  
EM260234.JPGJPG EM260234.JPG manage 208 K 29 Jan 2017 - 14:43 AlexHamilton  
EM260398.JPGJPG EM260398.JPG manage 216 K 09 Feb 2017 - 03:45 AlexHamilton  
EM260408.JPGJPG EM260408.JPG manage 220 K 09 Feb 2017 - 03:47 AlexHamilton  
EM260489.JPGJPG EM260489.JPG manage 95 K 09 Feb 2017 - 03:50 AlexHamilton  
EM260497.JPGJPG EM260497.JPG manage 192 K 09 Feb 2017 - 03:49 AlexHamilton  
EM260501.JPGJPG EM260501.JPG manage 211 K 09 Feb 2017 - 03:46 AlexHamilton  
EM260521.JPGJPG EM260521.JPG manage 207 K 09 Feb 2017 - 03:51 AlexHamilton  
EM270262.JPGJPG EM270262.JPG manage 232 K 29 Jan 2017 - 14:43 AlexHamilton  
IMG_5702_800x600.jpgjpg IMG_5702_800x600.jpg manage 102 K 24 Jan 2017 - 01:30 AlexHamilton  
IMG_5714.JPGJPG IMG_5714.JPG manage 128 K 29 Jan 2017 - 14:43 AlexHamilton  
P1130752-med.jpgjpg P1130752-med.jpg manage 209 K 26 Jan 2017 - 11:09 AlexHamilton  
P1130754-med.jpgjpg P1130754-med.jpg manage 250 K 26 Jan 2017 - 11:09 AlexHamilton  
Temperature_vs_time_31st.PNGPNG Temperature_vs_time_31st.PNG manage 33 K 09 Feb 2017 - 04:36 AlexHamilton  
Topic revision: r8 - 09 Feb 2017, AlexHamilton
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback